# IP-LLM: Semantic Space Informed Prompt Learning with LLM for Time Series Forecasting

### v1

@@ A.1 Implementation
We mainly follow the experimental configurations in (Wu et al., 2023) across all baselines within a unified evaluation pipeline, available at https://github.com/thuml/Time-Series-Library, for a fair comparison. We use GPT2-small (Radford et al., 2019) with the first 6 hidden layers enabled as the default backbone model. All our experiments are repeated three times and we report the averaged results. We implemented the model on PyTorch (Paszke et al., 2019) with all experiments conducted on NVIDIA RTX A6000 GPUs. We configure the patch length P as 16 with a stride S of 8. We maintain the number of semantic anchor candidates as 1000.
@@ A.2 Baseline Introduction
Time-LLM (Jin et al., 2024): Time-LLM reprograms time series tokens with NLP representation using multi-head attention and fine-tunes the pre-trained LLM with the prefix prompting to perform time series analysis.
@@ Appendix B Long-term Forecasting Results
Table 7 shows the detailed results of all prediction lengths of five Transformer-based forecasting models. IP-LLM shows a strong performance across different datasets compared to other transformer-based models. We also report the standard deviation of IP-LLM performance under three runs with different random seeds in Table 8, which exhibits that the performance of IP-LLM is relatively stable.
@@
| Methods | IP-LLM | FEDformer | Autoformer | Stationary | ETSformer | LightTS | |||||||
| MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | ||
| Weather | 96 | 0.050 | 0.121 | 0.217 | 0.296 | 0.266 | 0.336 | 0.173 | 0.223 | 0.197 | 0.281 | 0.182 | 0.242 |
| 192 | 0.090 | 0.164 | 0.276 | 0.336 | 0.307 | 0.367 | 0.245 | 0.285 | 0.237 | 0.312 | 0.227 | 0.287 | |
| 336 | 0.176 | 0.234 | 0.339 | 0.380 | 0.359 | 0.395 | 0.321 | 0.338 | 0.298 | 0.353 | 0.282 | 0.334 | |
| 720 | 0.300 | 0.306 | 0.403 | 0.428 | 0.419 | 0.428 | 0.414 | 0.410 | 0.352 | 0.288 | 0.352 | 0.386 | |
| Avg | 0.154 | 0.206 | 0.309 | 0.360 | 0.338 | 0.382 | 0.288 | 0.314 | 0.271 | 0.334 | 0.261 | 0.312 | |
| Electricity | 96 | 0.136 | 0.225 | 0.193 | 0.308 | 0.201 | 0.317 | 0.169 | 0.273 | 0.187 | 0.304 | 0.207 | 0.307 |
| 192 | 0.147 | 0.240 | 0.201 | 0.315 | 0.222 | 0.334 | 0.182 | 0.286 | 0.199 | 0.315 | 0.213 | 0.316 | |
| 336 | 0.155 | 0.243 | 0.214 | 0.329 | 0.231 | 0.338 | 0.200 | 0.304 | 0.212 | 0.329 | 0.230 | 0.333 | |
| 720 | 0.183 | 0.282 | 0.246 | 0.355 | 0.254 | 0.361 | 0.222 | 0.321 | 0.233 | 0.345 | 0.265 | 0.360 | |
| Avg | 0.155 | 0.247 | 0.214 | 0.327 | 0.227 | 0.338 | 0.193 | 0.296 | 0.208 | 0.323 | 0.229 | 0.329 | |
| Traffic | 96 | 0.347 | 0.231 | 0.587 | 0.366 | 0.613 | 0.388 | 0.612 | 0.338 | 0.607 | 0.392 | 0.615 | 0.391 |
| 192 | 0.353 | 0.240 | 0.604 | 0.373 | 0.616 | 0.382 | 0.613 | 0.340 | 0.621 | 0.399 | 0.601 | 0.382 | |
| 336 | 0.361 | 0.247 | 0.621 | 0.383 | 0.622 | 0.337 | 0.618 | 0.328 | 0.622 | 0.396 | 0.613 | 0.386 | |
| 720 | 0.392 | 0.278 | 0.626 | 0.382 | 0.660 | 0.408 | 0.653 | 0.355 | 0.632 | 0.396 | 0.658 | 0.407 | |
| Avg | 0.363 | 0.249 | 0.610 | 0.376 | 0.628 | 0.379 | 0.624 | 0.340 | 0.621 | 0.396 | 0.622 | 0.392 | |
| ETTh1 | 96 | 0.318 | 0.377 | 0.376 | 0.419 | 0.530 | 0.517 | 0.513 | 0.491 | 0.644 | 0.589 | 0.440 | 0.450 |
| 192 | 0.389 | 0.403 | 0.420 | 0.448 | 0.537 | 0.521 | 0.534 | 0.504 | 0.736 | 0.648 | 0.498 | 0.479 | |
| 336 | 0.427 | 0.423 | 0.459 | 0.465 | 0.596 | 0.583 | 0.588 | 0.535 | 0.827 | 0.707 | 0.550 | 0.510 | |
| 720 | 0.447 | 0.458 | 0.506 | 0.507 | 0.713 | 0.639 | 0.643 | 0.616 | 0.946 | 0.766 | 0.615 | 0.571 | |
| Avg | 0.396 | 0.416 | 0.440 | 0.460 | 0.594 | 0.565 | 0.570 | 0.537 | 0.788 | 0.677 | 0.526 | 0.502 | |
| ETTh2 | 96 | 0.245 | 0.313 | 0.358 | 0.397 | 0.454 | 0.490 | 0.476 | 0.458 | 0.340 | 0.391 | 0.408 | 0.445 |
| 192 | 0.324 | 0.365 | 0.429 | 0.439 | 0.486 | 0.517 | 0.512 | 0.493 | 0.430 | 0.439 | 0.561 | 0.526 | |
| 336 | 0.351 | 0.403 | 0.496 | 0.487 | 0.493 | 0.533 | 0.552 | 0.551 | 0.485 | 0.479 | 0.673 | 0.580 | |
| 720 | 0.363 | 0.410 | 0.463 | 0.474 | 0.515 | 0.543 | 0.562 | 0.560 | 0.500 | 0.497 | 1.006 | 0.721 | |
| Avg | 0.320 | 0.372 | 0.437 | 0.449 | 0.487 | 0.520 | 0.526 | 0.516 | 0.439 | 0.452 | 0.662 | 0.568 | |
| ETTm1 | 96 | 0.096 | 0.211 | 0.379 | 0.419 | 0.568 | 0.516 | 0.386 | 0.398 | 0.375 | 0.398 | 0.383 | 0.409 |
| 192 | 0.186 | 0.287 | 0.426 | 0.441 | 0.573 | 0.528 | 0.459 | 0.444 | 0.408 | 0.410 | 0.421 | 0.431 | |
| 336 | 0.286 | 0.352 | 0.445 | 0.459 | 0.587 | 0.534 | 0.495 | 0.464 | 0.435 | 0.428 | 0.454 | 0.456 | |
| 720 | 0.377 | 0.411 | 0.543 | 0.490 | 0.589 | 0.536 | 0.585 | 0.516 | 0.499 | 0.462 | 0.549 | 0.520 | |
| Avg | 0.236 | 0.315 | 0.448 | 0.452 | 0.579 | 0.529 | 0.481 | 0.456 | 0.429 | 0.425 | 0.452 | 0.454 | |
| ETTm2 | 96 | 0.082 | 0.194 | 0.203 | 0.287 | 0.287 | 0.359 | 0.192 | 0.274 | 0.189 | 0.280 | 0.239 | 0.335 |
| 192 | 0.163 | 0.263 | 0.269 | 0.328 | 0.325 | 0.388 | 0.280 | 0.339 | 0.253 | 0.319 | 0.346 | 0.412 | |
| 336 | 0.234 | 0.314 | 0.325 | 0.366 | 0.498 | 0.491 | 0.334 | 0.361 | 0.314 | 0.357 | 0.506 | 0.506 | |
| 720 | 0.337 | 0.362 | 0.421 | 0.415 | 0.548 | 0.517 | 0.417 | 0.413 | 0.414 | 0.413 | 0.702 | 0.606 | |
| Avg | 0.204 | 0.283 | 0.305 | 0.349 | 0.414 | 0.439 | 0.306 | 0.347 | 0.293 | 0.342 | 0.448 | 0.465 | |
| Dataset | Weather | Electricity | Traffic | ETTh1 | ETTh2 | ETTm1 | ETTm2 | |||||||
| Horizon | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE |
| 96 | 0.003 | 0.002 | 0.003 | 0.002 | 0.005 | 0.003 | 0.005 | 0.004 | 0.004 | 0.006 | 0.004 | 0.003 | 0.002 | 0.001 |
| 192 | 0.003 | 0.002 | 0.004 | 0.004 | 0.008 | 0.006 | 0.004 | 0.008 | 0.005 | 0.004 | 0.002 | 0.002 | 0.005 | 0.005 |
| 336 | 0.003 | 0.003 | 0.004 | 0.005 | 0.005 | 0.004 | 0.005 | 0.004 | 0.006 | 0.007 | 0.007 | 0.005 | 0.002 | 0.003 |
| 720 | 0.005 | 0.002 | 0.003 | 0.003 | 0.006 | 0.004 | 0.004 | 0.003 | 0.005 | 0.005 | 0.004 | 0.003 | 0.003 | 0.004 |
@@ Appendix C Full Short-term Forecasting Results
Table 9 shows the full short-term forecasting experiment results on M4 datasets. IP-LLM consistently outperforms the majority of baseline models in most cases. It surpasses the performance of OFA significantly and achieves comparable forecasting performance compared to Time-LLM, which can be attributed to proposed semantic space informed prompting.
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | N-HiTS | N-BEATS | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | |
| Year. | SMAPE | 13.271 | 13.419 | 15.11 | 13.652 | 16.965 | 13.477 | 13.422 | 13.487 | 15.378 | 14.021 | 13.974 | 14.727 | 18.009 |
| MASE | 2.973 | 3.005 | 3.565 | 3.095 | 4.283 | 3.019 | 3.056 | 3.036 | 3.554 | 3.036 | 3.134 | 3.078 | 4.487 | |
| OWA | 0.781 | 0.789 | 0.911 | 0.807 | 1.058 | 0.792 | 0.795 | 0.795 | 0.918 | 0.811 | 0.822 | 0.807 | 1.115 | |
| Quart. | SMAPE | 10.032 | 10.110 | 10.597 | 10.353 | 12.145 | 10.38 | 10.185 | 10.564 | 10.465 | 11.100 | 11.338 | 10.958 | 13.376 |
| MASE | 1.175 | 1.178 | 1.253 | 1.209 | 1.520 | 1.233 | 1.18 | 1.252 | 1.227 | 1.35 | 1.365 | 1.325 | 1.906 | |
| OWA | 0.883 | 0.889 | 0.938 | 0.911 | 1.106 | 0.921 | 0.893 | 0.936 | 0.923 | 0.996 | 1.012 | 0.981 | 1.302 | |
| Month. | SMAPE | 12.993 | 12.980 | 13.258 | 13.079 | 13.514 | 12.959 | 13.059 | 13.089 | 13.513 | 14.403 | 13.958 | 13.917 | 14.588 |
| MASE | 0.969 | 0.963 | 1.003 | 0.974 | 1.037 | 0.970 | 1.013 | 0.996 | 1.039 | 1.147 | 1.103 | 1.097 | 1.368 | |
| OWA | 0.909 | 0.903 | 0.931 | 0.911 | 0.956 | 0.905 | 0.929 | 0.922 | 0.957 | 1.038 | 1.002 | 0.998 | 1.149 | |
| Others. | SMAPE | 4.803 | 4.795 | 6.124 | 4.78 | 6.709 | 4.952 | 4.711 | 6.599 | 6.913 | 7.148 | 5.485 | 6.302 | 7.267 |
| MASE | 3.241 | 3.178 | 4.116 | 3.231 | 4.953 | 3.347 | 3.054 | 4.430 | 4.507 | 4.064 | 3.865 | 4.064 | 5.240 | |
| OWA | 1.017 | 1.006 | 1.259 | 1.012 | 1.487 | 1.049 | 0.977 | 1.393 | 1.438 | 1.304 | 1.187 | 1.304 | 1.591 | |
| Avg. | SMAPE | 11.832 | 11.983 | 12.690 | 12.142 | 13.639 | 12.059 | 12.035 | 12.250 | 12.880 | 13.160 | 12.909 | 12.780 | 14.718 |
| MASE | 1.584 | 1.595 | 1.808 | 1.631 | 2.095 | 1.623 | 1.625 | 1.698 | 1.836 | 1.775 | 1.771 | 1.756 | 2.408 | |
| OWA | 0.830 | 0.859 | 0.94 | 0.874 | 1.051 | 0.869 | 0.869 | 0.896 | 0.955 | 0.949 | 0.939 | 0.930 | 1.172 | |

### v2

@@ A.1 Implementation
We mainly follow the experimental configurations in (Wu et al., 2023) across all baselines within a unified evaluation pipeline, available at https://github.com/thuml/Time-Series-Library, for a fair comparison. We use GPT2-small (Radford et al., 2019) with the first 6 hidden layers enabled as the default backbone model. All our experiments are repeated three times and we report the averaged results. We implemented the model on PyTorch (Paszke et al., 2019) with all experiments conducted on NVIDIA RTX A6000 GPUs and NVIDIA A100 GPUs. We configure the patch length P as 16 with a stride S of 8. Our experimental results were obtained by performing a grid search over combinations within the following search spaces: long-term trend length (24, 48, 96, 144), seasonal trend length (2, 4, 8, 12, 24), prompt length (2, 4, 8, 16, 32), and semantic space size (1000, 2000, 5000). Our code is available at https://github.com/panzijie825/S2IP-LLM.
@@ A.2 Baseline Introduction
Time-LLM (Jin et al., 2024): Time-LLM reprograms time series tokens with NLP representation using multi-head attention and fine-tunes the pre-trained LLM with the prefix prompting to perform time series analysis. We reproduced and reported the experimental results for long-term forecasting using two LLM backbones: ‘L’ represents LLaMA (Touvron et al., 2023a), and ‘G’ represents GPT-2 (Radford et al., 2019).
@@ Appendix B Long-term Forecasting Results
Table 6 shows the detailed results of all prediction lengths of five Transformer-based forecasting models. IP-LLM shows a strong and relatively stable performance across different datasets compared to other transformer-based models.
@@
| Methods | IP-LLM | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | LightTS | ||||||||
| MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | ||
| Weather | 96 | 0.145 | 0.195 | 0.172 | 0.220 | 0.217 | 0.296 | 0.266 | 0.336 | 0.173 | 0.223 | 0.197 | 0.281 | 0.182 | 0.242 |
| 192 | 0.190 | 0.235 | 0.219 | 0.261 | 0.276 | 0.336 | 0.307 | 0.367 | 0.245 | 0.285 | 0.237 | 0.312 | 0.227 | 0.287 | |
| 336 | 0.243 | 0.280 | 0.280 | 0.306 | 0.339 | 0.380 | 0.359 | 0.395 | 0.321 | 0.338 | 0.298 | 0.353 | 0.282 | 0.334 | |
| 720 | 0.312 | 0.326 | 0.365 | 0.359 | 0.403 | 0.428 | 0.419 | 0.428 | 0.414 | 0.410 | 0.352 | 0.288 | 0.352 | 0.386 | |
| Avg | 0.222 | 0.259 | 0.259 | 0.287 | 0.309 | 0.360 | 0.338 | 0.382 | 0.288 | 0.314 | 0.271 | 0.334 | 0.261 | 0.312 | |
| Electricity | 96 | 0.135 | 0.230 | 0.168 | 0.272 | 0.193 | 0.308 | 0.201 | 0.317 | 0.169 | 0.273 | 0.187 | 0.304 | 0.207 | 0.307 |
| 192 | 0.149 | 0.247 | 0.184 | 0.289 | 0.201 | 0.315 | 0.222 | 0.334 | 0.182 | 0.286 | 0.199 | 0.315 | 0.213 | 0.316 | |
| 336 | 0.167 | 0.266 | 0.198 | 0.300 | 0.214 | 0.329 | 0.231 | 0.338 | 0.200 | 0.304 | 0.212 | 0.329 | 0.230 | 0.333 | |
| 720 | 0.200 | 0.287 | 0.220 | 0.320 | 0.246 | 0.355 | 0.254 | 0.361 | 0.222 | 0.321 | 0.233 | 0.345 | 0.265 | 0.360 | |
| Avg | 0.161 | 0.257 | 0.192 | 0.295 | 0.214 | 0.327 | 0.227 | 0.338 | 0.193 | 0.296 | 0.208 | 0.323 | 0.229 | 0.329 | |
| Traffic | 96 | 0.379 | 0.274 | 0.593 | 0.321 | 0.587 | 0.366 | 0.613 | 0.388 | 0.612 | 0.338 | 0.607 | 0.392 | 0.615 | 0.391 |
| 192 | 0.397 | 0.282 | 0.617 | 0.336 | 0.604 | 0.373 | 0.616 | 0.382 | 0.613 | 0.340 | 0.621 | 0.399 | 0.601 | 0.382 | |
| 336 | 0.407 | 0.289 | 0.629 | 0.336 | 0.621 | 0.383 | 0.622 | 0.337 | 0.618 | 0.328 | 0.622 | 0.396 | 0.613 | 0.386 | |
| 720 | 0.440 | 0.301 | 0.640 | 0.350 | 0.626 | 0.382 | 0.660 | 0.408 | 0.653 | 0.355 | 0.632 | 0.396 | 0.658 | 0.407 | |
| Avg | 0.405 | 0.286 | 0.620 | 0.336 | 0.610 | 0.376 | 0.628 | 0.379 | 0.624 | 0.340 | 0.621 | 0.396 | 0.622 | 0.392 | |
| ETTh1 | 96 | 0.366 | 0.396 | 0.468 | 0.475 | 0.376 | 0.419 | 0.530 | 0.517 | 0.513 | 0.491 | 0.644 | 0.589 | 0.440 | 0.450 |
| 192 | 0.401 | 0.420 | 0.484 | 0.485 | 0.420 | 0.448 | 0.537 | 0.521 | 0.534 | 0.504 | 0.736 | 0.648 | 0.498 | 0.479 | |
| 336 | 0.412 | 0.431 | 0.536 | 0.516 | 0.459 | 0.465 | 0.596 | 0.583 | 0.588 | 0.535 | 0.827 | 0.707 | 0.550 | 0.510 | |
| 720 | 0.440 | 0.458 | 0.593 | 0.537 | 0.506 | 0.507 | 0.713 | 0.639 | 0.643 | 0.616 | 0.946 | 0.766 | 0.615 | 0.571 | |
| Avg | 0.406 | 0.427 | 0.520 | 0.505 | 0.440 | 0.460 | 0.594 | 0.565 | 0.570 | 0.537 | 0.788 | 0.677 | 0.526 | 0.502 | |
| ETTh2 | 96 | 0.278 | 0.340 | 0.376 | 0.415 | 0.358 | 0.397 | 0.454 | 0.490 | 0.476 | 0.458 | 0.340 | 0.391 | 0.408 | 0.445 |
| 192 | 0.346 | 0.385 | 0.409 | 0.440 | 0.429 | 0.439 | 0.486 | 0.517 | 0.512 | 0.493 | 0.430 | 0.439 | 0.561 | 0.526 | |
| 336 | 0.367 | 0.406 | 0.425 | 0.455 | 0.496 | 0.487 | 0.493 | 0.533 | 0.552 | 0.551 | 0.485 | 0.479 | 0.673 | 0.580 | |
| 720 | 0.400 | 0.436 | 0.488 | 0.494 | 0.463 | 0.474 | 0.515 | 0.543 | 0.562 | 0.560 | 0.500 | 0.497 | 1.006 | 0.721 | |
| Avg | 0.347 | 0.391 | 0.425 | 0.451 | 0.437 | 0.449 | 0.487 | 0.520 | 0.526 | 0.516 | 0.439 | 0.452 | 0.662 | 0.568 | |
| ETTm1 | 96 | 0.288 | 0.346 | 0.329 | 0.377 | 0.379 | 0.419 | 0.568 | 0.516 | 0.386 | 0.398 | 0.375 | 0.398 | 0.383 | 0.409 |
| 192 | 0.323 | 0.365 | 0.371 | 0.401 | 0.426 | 0.441 | 0.573 | 0.528 | 0.459 | 0.444 | 0.408 | 0.410 | 0.421 | 0.431 | |
| 336 | 0.359 | 0.390 | 0.417 | 0.428 | 0.445 | 0.459 | 0.587 | 0.534 | 0.495 | 0.464 | 0.435 | 0.428 | 0.454 | 0.456 | |
| 720 | 0.403 | 0.418 | 0.483 | 0.464 | 0.543 | 0.490 | 0.589 | 0.536 | 0.585 | 0.516 | 0.499 | 0.462 | 0.549 | 0.520 | |
| Avg | 0.343 | 0.379 | 0.400 | 0.417 | 0.448 | 0.452 | 0.579 | 0.529 | 0.481 | 0.456 | 0.429 | 0.425 | 0.452 | 0.454 | |
| ETTm2 | 96 | 0.165 | 0.257 | 0.201 | 0.286 | 0.203 | 0.287 | 0.287 | 0.359 | 0.192 | 0.274 | 0.189 | 0.280 | 0.239 | 0.335 |
| 192 | 0.222 | 0.299 | 0.260 | 0.329 | 0.269 | 0.328 | 0.325 | 0.388 | 0.280 | 0.339 | 0.253 | 0.319 | 0.346 | 0.412 | |
| 336 | 0.277 | 0.330 | 0.331 | 0.376 | 0.325 | 0.366 | 0.498 | 0.491 | 0.334 | 0.361 | 0.314 | 0.357 | 0.506 | 0.506 | |
| 720 | 0.363 | 0.390 | 0.428 | 0.430 | 0.421 | 0.415 | 0.548 | 0.517 | 0.417 | 0.413 | 0.414 | 0.413 | 0.702 | 0.606 | |
| Avg | 0.257 | 0.319 | 0.305 | 0.355 | 0.305 | 0.349 | 0.414 | 0.439 | 0.306 | 0.347 | 0.293 | 0.342 | 0.448 | 0.465 | |
@@ Appendix C Full Short-term Forecasting Results
Table 7 shows the full short-term forecasting experiment results on M4 datasets. IP-LLM consistently outperforms the majority of baseline models in most cases. It surpasses the performance of OFA significantly and achieves slightly better forecasting performance than PatchTST, which can be attributed to proposed semantic space informed prompting.
| Methods | IP-LLM | Time-LLM(G) | OFA | iTrans. | Dlinear | PatchTST | N-HiTS | N-BEATS | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | |
| Year. | SMAPE | 13.413 | 13.75 | 15.11 | 13.652 | 16.965 | 13.477 | 13.422 | 13.487 | 15.378 | 14.021 | 13.974 | 14.727 | 18.009 |
| MASE | 3.024 | 3.055 | 3.565 | 3.095 | 4.283 | 3.019 | 3.056 | 3.036 | 3.554 | 3.036 | 3.134 | 3.078 | 4.487 | |
| OWA | 0.792 | 0.805 | 0.911 | 0.807 | 1.058 | 0.792 | 0.795 | 0.795 | 0.918 | 0.811 | 0.822 | 0.807 | 1.115 | |
| Quart. | SMAPE | 10.352 | 10.671 | 10.597 | 10.353 | 12.145 | 10.38 | 10.185 | 10.564 | 10.465 | 11.100 | 11.338 | 10.958 | 13.376 |
| MASE | 1.228 | 1.276 | 1.253 | 1.209 | 1.520 | 1.233 | 1.18 | 1.252 | 1.227 | 1.35 | 1.365 | 1.325 | 1.906 | |
| OWA | 0.922 | 0.95 | 0.938 | 0.911 | 1.106 | 0.921 | 0.893 | 0.936 | 0.923 | 0.996 | 1.012 | 0.981 | 1.302 | |
| Month. | SMAPE | 12.995 | 13.416 | 13.258 | 13.079 | 13.514 | 12.959 | 13.059 | 13.089 | 13.513 | 14.403 | 13.958 | 13.917 | 14.588 |
| MASE | 0.97 | 1.045 | 1.003 | 0.974 | 1.037 | 0.970 | 1.013 | 0.996 | 1.039 | 1.147 | 1.103 | 1.097 | 1.368 | |
| OWA | 0.91 | 0.957 | 0.931 | 0.911 | 0.956 | 0.905 | 0.929 | 0.922 | 0.957 | 1.038 | 1.002 | 0.998 | 1.149 | |
| Others. | SMAPE | 4.805 | 4.973 | 6.124 | 4.78 | 6.709 | 4.952 | 4.711 | 6.599 | 6.913 | 7.148 | 5.485 | 6.302 | 7.267 |
| MASE | 3.247 | 3.412 | 4.116 | 3.231 | 4.953 | 3.347 | 3.054 | 4.430 | 4.507 | 4.064 | 3.865 | 4.064 | 5.240 | |
| OWA | 1.017 | 1.053 | 1.259 | 1.012 | 1.487 | 1.049 | 0.977 | 1.393 | 1.438 | 1.304 | 1.187 | 1.304 | 1.591 | |
| Avg. | SMAPE | 12.021 | 12.494 | 12.690 | 12.142 | 13.639 | 12.059 | 12.035 | 12.250 | 12.880 | 13.160 | 12.909 | 12.780 | 14.718 |
| MASE | 1.612 | 1.731 | 1.808 | 1.631 | 2.095 | 1.623 | 1.625 | 1.698 | 1.836 | 1.775 | 1.771 | 1.756 | 2.408 | |
| OWA | 0.857 | 0.913 | 0.94 | 0.874 | 1.051 | 0.869 | 0.869 | 0.896 | 0.955 | 0.949 | 0.939 | 0.930 | 1.172 | |